# MAME: Multidimensional Adaptive Metamer Exploration with Human Perceptual Feedback

### v1

@@ Abstract
Alignment between human brain networks and artificial models has become an active research area in both machine learning and neuroscience. A widely adopted approach to explore the functional alignment between them is to identify “metamers” for both humans and models. Metamers refer to input stimuli that are physically different but equivalent within a given system. If a model’s metameric space completely matched the human metameric space, the model would achieve functional alignment with humans. However, conventional methods lack a direct approach to searching for the human metameric space. As a result, researchers have had to first develop biologically inspired models and then infer about human metamers indirectly by testing whether model metamers also appear as metamers to humans. Here, we propose the Multidimensional Adaptive Metamer Exploration (MAME) framework, which enables direct high-dimensional exploration of human metameric space. MAME leverages online image generation guided by human perceptual feedback, allowing for a more flexible search process that mitigates the exploration constraint of human metamers. Specifically, the MAME framework modulates reference images across multiple dimensions by leveraging hierarchical responses from convolutional neural networks (CNNs). The generated metameric images are presented to human participants, and their perceptual discriminability from the reference images is assessed through a behavioral task. Based on participants’ responses, subsequent image generation parameters are adaptively updated online. Using our MAME framework, we successfully measured a human metameric space of over fifty dimensions within a single experiment. Experimental results showed that human discrimination sensitivity was lower for metameric images based on low-level features compared to high-level features, which image contrast metrics could not explain. The finding suggests that the model computes low-level information that is not essential for human perception. Our MAME framework has the potential to contribute to developing interpretable AI and to advance our understanding of brain function in neuroscience.
Keywords: cognitive neuroscience; vision science; model metamers; texture synthesis; psychophysics
@@ 1.1 Functional Alignment Between Human Vision and Artificial Neural Networks
Recent advancements in artificial neural networks (ANNs) have not only improved their ability to solve visual tasks but have also positioned them as valuable analogies for understanding human visual processes. In particular, deep convolutional neural networks (CNNs) have been noted for their structural and functional similarities to the hierarchical processing observed in biological vision. These models have demonstrated considerable success in approximating the information flow from the primary visual cortex (V1) to higher visual areas such as V4 and the inferotemporal (IT) cortex in the ventral visual stream (Yamins and DiCarlo,, 2016). However, numerous studies have reported significant divergences between the behavioral responses of humans and ANN models (Geirhos et al., 2018a, ; Geirhos et al., 2018b, ; Szegedy,, 2013). Consequently, extensive research efforts have been directed toward achieving a functional alignment between human vision and artificial models (Geirhos et al.,, 2021).
1.2 Metamers as a tool for functional understanding of cognitive neural process
@@
Through this metamer-based approach, various visual functions have been uncovered in the history of vision sciences. A classic example is the finding that human color vision is mediated by three types of sensors (reviewed by Gegenfurtner, (2003)). The discovery that humans do not directly sense light as a function of wavelength was based on experimental results showing that physically different spectral compositions of light can appear identical to human observers, i.e., they form metamers.
Figure 1: The MAME framework assumes that images can be generated at arbitrary locations starting from a reference stimulus.
@@
The approach of using metamers for functional understanding has further advanced with the use of deep neural network models as artificial systems. Gatys et al., (2015) proposed an image generation technique that utilizes intermediate features in convolutional neural networks, summarizing them with statistical representations such as Gram matrices. This method has been particularly beneficial for cognitive neuroscience researchers interested in higher-order hierarchical visual processing. For example, in the context of object recognition research, a metamer-based approach combining this image algorithm with fMRI measurements has demonstrated that object-selective regions in higher ventral visual areas exhibit texture-like representation, computed by Gram matrices (Jagadeesh and Gardner,, 2022).
When exploring human metamers using image generation algorithms from deep learning models, one first generates model metamers from noise images for a specific target image and then tests whether humans perceive the generated images as equivalent to the target (Figure 1, left). In this case, the model used for metamer generation should closely align with human vision. In fact, adversarially-trained CNNs (Salman et al.,, 2020), which are known as robust for various tasks like humans, can generate metamers that better explain human peripheral vision (Harrington and Deza,, 2021). Furthermore, there have been attempts to evaluate the alignment between humans and models by assessing the extent to which metamers generated by models align with human metamers (Feather et al.,, 2019, 2023).
@@ 1.3 Challenges in Exploring High-Dimensional Human Metameric Spaces
A key limitation of the conventional metamer approach (Figure 1, left) is that the exploration of human metamers is indirect. First, a biologically plausible model must be prepared, followed by the generation of model metamers. Then, it is necessary to test whether these model metamers are also perceived as metamers by humans. This method can completely reveal the structure of the human metameric space only if a model that already aligns with human perception is available beforehand.
@@
Figure 2: The MAME framework creates tasks adaptively online and efficiently explores metamers.
  • It can explore more than 50 dimensions of human metameric space simultaneously in one psychophysical experiment.
@@
  • The experimental environment is highly flexible because online image generation is performed on a Python server using a GPU, while its human experiment is conducted via a web browser with JavaScript. Specifically, it can be easily applied to crowdsourcing or fMRI experiments that need online feedback.
1.4 Proposed Approach: The MAME Framework
In this study, we propose a framework for exploring high-dimensional human metameric spaces directly, called Multidimensional Adaptive Metamer Exploration (MAME) (Figure 2). The MAME framework enables direct multidimensional exploration of the human metameric space by utilizing online image generation based on human feedback, thereby relaxing the constraints that a human-aligned model is needed beforehand. Specifically, in the MAME framework, the model is used solely to determine the search direction for human metamers Figure 1 (right). While it is better for the model to be biologically plausible when determining the direction, the final human metameric boundary can be determined independent of the model’s metamers. This aspect relaxes the constraints on human metamer exploration. Furthermore, to enhance the efficiency of multidimensional exploration, we employ online image generation guided by trial-wise human feedback, enabling the adaptive search for metameric boundaries.
@@ 2.1 The MAME Framework
The MAME framework is designed to adaptively explore the boundaries of the human metameric space. This exploration is achieved by systematically perturbing a given reference image in multiple dimensions Fig. (1, right). The execution of the MAME framework relies on the following requirements:
@@
    Feedback from Tasks performed by Humans or Animals Finally, our MAME framework requires feedback from tasks performed by humans or animals. As described in the Introduction, metamers refer to images that are physically different but equivalent within a given system. However, the definition of “equivalence” changes across studies. Some adopt categorical equivalence based on behavioral responses (Feather et al.,, 2019), while others define it in terms of perceptual appearance (Freeman and Simoncelli,, 2011; Deza et al.,, 2017) using ABX tasks. In general, measuring perceptual appearance psychophysically is more time-consuming than categorical judgment. Moreover, evaluating the equivalence of neural or BOLD activations in neurophysiological or fMRI studies requires substantial costs. Our method can be applied to feedback from any type of task; however, it is particularly effective for tasks where response acquisition takes longer, as online image generation allows for a more efficient increase in the number of trials.
In the present study, the MAME framework was implemented by (1) defining the direction of human metameric space using a CNN and ICA, (2) implementing a stimulus generation method that perturbs the reference stimulus via gradient descent, and (3) incorporating ABX tasks as feedback.
@@ 2.2 Definition of the exploration direction of human metameric space
The definition of the metameric space used in this study is illustrated in Fig. 3. This definition is based on a CNN and independent component analysis (ICA). First, natural images were passed through a pre-trained CNN, while extracting Gram matrices from the activations of each layer as position-invariant features. Here, we used Gram matrices because the present study explored metamers based on the equivalence of perceptual appearance and focused summary statistics like textures representations (Jagadeesh and Gardner,, 2022). For the model, we used a robust variant of ResNet50 (He et al.,, 2016) (ResNet50 robustness ImageNet L2-Norm $\epsilon=3.0$) (Engstrom et al.,, 2019). This model was trained to be resistant to adversarial attacks and is known for having internal representations that are biologically plausible (Feather et al.,, 2023; Gaziv et al.,, 2024).
Next, ICA was performed on these Gram matrices, and components with high explanatory power were selected, thereby considering images as points in the ICA coordinate space. The exploration direction was defined along the axes of ICA from the reference stimulus.
Figure 3: Metameric space was defined using a CNN and ICA.
In the experiment, the metameric space was explored in a 54-dimensional space, determined as follows. Feature extraction layers in the CNN included conv1, layer3, and avgpool to explore low-, middle-, and high-level features. For ICA, the top three components from each layer were selected. Each component was explored in both positive and negative directions. Additionally, three conditions for stimulus presentation eccentricities (4°, 8°, and 12°) were considered. Consequently, the total dimensionality was calculated as 54 dimensions (3 x 3 x 3 x 2).
2.2.1 Image Dataset
The ImageNet validation dataset (Deng et al.,, 2009) was used as a set of natural images. Each image was first resized to 256×256 pixels, and then the central 224×224 pixels were cropped for use to match the input image set and image size used during the training of the pre-trained model.
2.2.2 Feature Extraction
@@
Gram matrices were extracted from each layer of the robust ResNet50 model as features of the input images. A Gram matrix represents the similarity of activations across multiple filters, summed over spatial positions. Given the activation of the $l$-th layer of a CNN for an input image as $F_{jk}$, where $j$ denotes the filter index and $k$ denotes the spatial position, the Gram matrix $G_{ij}$ is defined as:
$G_{ij}=\sum_{k}F_{ik}F_{jk}.$
@@
In this study, Gram matrices were extracted from multiple layers of the model using 1,000 images randomly selected from the ImageNet validation dataset. Since Gram matrices discard spatial information, they are well-suited for extracting spatially invariant features of images (Gatys et al.,, 2015, 2016).
2.2.3 Independent Component Analysis (ICA)
@@
ICA was employed in this study to map the gram matrix feature onto a coordinate space. ICA assumes non-Gaussianity of signals and computes a linear transformation $W$ via gradient descent to separate source signals $X$ into independent components $S$ as $S=WX$. Each independent component (hereafter referred to as an ICA component) is standardized to have zero mean and a covariance matrix equal to the identity matrix.
In this study, ICA was applied to the dataset as follows. The Gram matrices of the target layers were vectorized to form feature vectors $x$ for each image. Using the feature vector x of 1,000 images, an input data matrix $X$ was constructed, and ICA was applied to obtain the independent component matrix:
$S=WX,\quad\mathbf{s}=W\mathbf{x},\quad X=SA^{\mathsf{T}}$
@@
where $A$ represents the mixing matrix. The number of independent components was pre-set to 100, and the top 3 components with the highest contribution were selected. The contribution of each component was quantified using the explained variance (EV) defined as:
$\mathrm{EV}_{i}=1-\frac{\|X-\hat{X}_{i}\|_{\text{F}}^{2}}{\|X\|_{\text{F}}^{2}},$
@@
$\hat{X}_{i}=S_{i}A^{\mathsf{T}},$
with $S_{i}$ being the matrix retaining only the $i$-th independent component. Unlike PCA, the total explained variance for multiple ICA components does not necessarily match the sum of the individual component contributions. The contribution of each component is shown in Fig. 4.
Figure 4: Explained variance of ICA components sorted in descending order. (Top-left) conv1, (Top-right) layer1, (Middle-left) layer2, (Middle-right) layer3, (Bottom-left) layer4, (Bottom-right) avgpool.
2.2.4 Selection of the Reference Stimuli Used in the Experiment
In this study, we focused on metamers of images located near the zero point of the ICA components for low-, mid-, and high-level features within the natural image distribution. The reason for focusing on them is to sample many reference images for our behavioral experiment so that participants do not memorize individual images.
From 50,000 images in the ImageNet validation dataset, we sampled images for three target layers whose ICA component distances from zero were in the lowest 20%. Among these, 614 images that were selected across all three layers were used as experimental stimuli.

### v2

@@ Abstract
Alignment between human brain networks and artificial models has become an active research area in vision science and machine learning. A widely adopted approach is identifying "metamers," stimuli physically different yet perceptually equivalent within a system. However, conventional methods lack a direct approach to searching for the human metameric space. Instead, researchers first develop biologically inspired models and then infer about human metamers indirectly by testing whether model metamers also appear as metamers to humans. Here, we propose the Multidimensional Adaptive Metamer Exploration (MAME) framework, enabling direct, high-dimensional exploration of human metameric spaces through online image generation guided by human perceptual feedback. MAME modulates reference images across multiple dimensions based on hierarchical neural network responses, adaptively updating generation parameters according to participants’ perceptual discriminability. Using MAME, we successfully measured multidimensional human metameric spaces within a single psychophysical experiment. Experimental results using a biologically plausible CNN model showed that human discrimination sensitivity was lower for metameric images based on low-level features compared to high-level features, which image contrast metrics could not explain. The finding suggests a relatively worse alignment between the metameric spaces of humans and the CNN model for low-level processing compared to high-level processing. Counterintuitively, given recent discussions on alignment at higher representational levels, our results highlight the importance of early visual computations in shaping biologically plausible models. Our MAME framework can serve as a future scientific tool for directly investigating the functional organization of human vision.
Keywords: metamers; texture synthesis; psychophysics; cognitive neuroscience; machine learning
@@ 1.1 Functional Alignment Between Human Vision and Artificial Neural Networks
Recent advancements in artificial neural networks (ANNs) have not only improved their ability to solve visual tasks but have also positioned them as valuable analogies for understanding human visual processes. In particular, deep convolutional neural networks (CNNs) have been noted for their structural and functional similarities to the hierarchical processing observed in biological vision. These models have demonstrated considerable success in approximating the information flow from the primary visual cortex to higher visual areas in the ventral visual stream (Yamins and DiCarlo,, 2016). However, numerous studies have reported significant divergences between the behavioral responses of humans and ANN models (Geirhos et al., 2018a, ; Geirhos et al., 2018b, ; Szegedy,, 2013). Consequently, extensive research efforts have been directed toward achieving a functional alignment between human vision and artificial models (Geirhos et al.,, 2021).
1.2 Metamer exploration as a tool for functional understanding of cognitive neural processes
@@
Through this metamer-based approach, various visual functions have been uncovered in the history of vision sciences. A classic example is the finding that human color vision is mediated by three types of sensors (reviewed by (Gegenfurtner,, 2003)). The discovery that humans do not directly sense light as a function of wavelength was based on experimental results showing that physically different spectral compositions of light can appear identical to human observers, i.e., they form metamers.
@@
The approach of using metamers for functional understanding has further advanced with the use of deep neural network models (DNNs) as artificial systems. Gatys et al. (Gatys et al.,, 2015) proposed an image generation technique that utilizes intermediate features in convolutional neural networks, summarizing them with statistical representations such as Gram matrices. This method has been particularly beneficial for cognitive neuroscience researchers interested in higher-order hierarchical visual processing. For example, in the context of object recognition research, a metamer-based approach combining this image algorithm with fMRI measurements has demonstrated that object-selective regions in higher ventral visual areas exhibit texture-like representation, computed by Gram matrices (Jagadeesh and Gardner,, 2022). Furthermore, there have been attempts to evaluate the alignment between humans and models by assessing the extent to which metamers generated by DNN models align with human metamers (Feather et al.,, 2019, 2023).
@@ 1.3 Challenges in Exploring High-Dimensional Human Metameric Spaces
A key limitation of approaches that use deep neural networks (DNNs) to explore human metamers is that these approaches are inherently indirect. This limitation becomes evident when examining how previous studies have explored human metamers (Figure 1).
Figure 1: A conceptual illustration of the MAME framework. In previous studies, human metamers were explored indirectly by generating model metamers from noise images and presenting them to human observers. In contrast, our approach proposes a method for directly measuring human metamers by performing multidimensional exploration starting from a reference image.
In previous studies, model metamers are first synthesized from noise to match a target image in a given DNN, and human observers are then asked whether these images appear equivalent to the original target (Figure 1, left). In this approach, the effectiveness of human metamer exploration critically depends on how well the model used for synthesis aligns with human visual perception. Indeed, adversarially trained convolutional neural networks (CNNs), which are known to exhibit human-like robustness across various tasks (Salman et al.,, 2020), have been shown to generate metamers that better account for human peripheral vision (Harrington and Deza,, 2021).
However, this method can fully reveal the structure of human metameric representations only if a suitably aligned model is available in advance. In this sense, human metamer exploration remains an indirect approach.
@@
1.4 Proposed Approach: The MAME Framework
In this study, we propose a framework for exploring high-dimensional human metameric spaces directly, called Multidimensional Adaptive Metamer Exploration (MAME) (Figure 2). The MAME framework enables direct multidimensional exploration of the human metameric space by utilizing online image generation based on human feedback, thereby relaxing the constraints that a human-aligned model is needed beforehand. Specifically, in the MAME framework, the model is used solely to determine the search direction for human metamers. While it is better for the model to be biologically plausible when determining the direction, the final human metameric boundary can be determined independent of the model’s metamers. This aspect relaxes the constraints on human metamer exploration. Furthermore, to enhance the efficiency of multidimensional exploration, we employ online image generation guided by trial-wise human feedback, enabling the adaptive search for metameric boundaries.
Figure 2: Exploring human metamers using the MAME framework.The exploration space is defined by the model (in this study, ICA components of CNN activations). Stimuli are synthesized along the defined directions, and their perceptual equivalence to a reference image is evaluated via a human behavioral task. Based on the feedback, the estimated human metamer is adaptively updated, and the task is repeated until convergence.
  • It can explore multidimensional human metameric space simultaneously in one psychophysical experiment.
@@ 1.4 Proposed Approach: The MAME Framework
  • The experimental environment is highly flexible because online image generation is performed on a Python server using a GPU, while the human experiment is conducted via a web browser with JavaScript.
@@ 2.1 The MAME Framework
The MAME framework is designed to adaptively explore the boundaries of the human metameric space. This exploration is achieved by systematically perturbing a given reference image in multiple dimensions (Fig. 1, right). The execution of the MAME framework relies on the following requirements:
@@
    Feedback from Tasks performed by Humans or Animals Finally, our MAME framework requires feedback from tasks performed by humans or animals. As described in the Introduction, metamers refer to images that are physically different but equivalent within a given system. However, the definition of “equivalence” changes across studies. Some adopt categorical equivalence based on behavioral responses (Feather et al.,, 2019), while others define it in terms of perceptual appearance (Freeman and Simoncelli,, 2011; Deza et al.,, 2017) using ABX tasks. Our method can be applied to feedback from any type of task if the computation speed of image generation is fast enough for real-time online experiments.
Depending on the research objective, different methods can be selected for each requirement. In the present study, our aim was to demonstrate the effectiveness of MAME, a novel comprehensive framework, under specific conditions. To this end, we evaluated its effectiveness using a canonical approach within the context of metamer exploration research. Specifically, the MAME framework was implemented by (1) defining the direction of human metameric space using a CNN and ICA, (2) implementing a stimulus generation method that perturbs the reference stimulus via gradient descent, and (3) incorporating ABX tasks as feedback. The specific motivations for selecting each method are described in detail in the corresponding sections of the Methods.
@@ 2.2 Definition of the exploration direction of human metameric space
In this section, we describe how the metameric search space was defined in the present study. We constructed the search space based on the activations of a robust variant of ResNet50 (He et al.,, 2016) (ResNet50 robustness ImageNet L2-Norm $\epsilon=3.0$) (Engstrom et al.,, 2019). Among the many available DNN models, we selected this specific model because previous work has shown it to best approximate human metamers under indirect exploration paradigms (Feather et al.,, 2023). In our MAME framework, the model is used to define the direction of exploration. Therefore, in order to efficiently estimate human metamers under a limited number of conditions, it is preferable that the model be biologically plausible. In particular, we note that manipulating high-level features is more difficult in other models, as shown in the comparison results provided in the Supplementary Material (Fig. S1).
@@
Next, we detail how the search space was concretely defined (Fig. 3). First, natural images were passed through the robust ResNet50, while extracting Gram matrices from the activations of each layer as position-invariant features. Here, we used Gram matrices because the present study explored metamers based on the equivalence of perceptual appearance and focused summary statistics like textures representations (Gatys et al.,, 2015, 2016; Jagadeesh and Gardner,, 2022).
Figure 3: Definition of exploration directions in the present study. An input image is passed through a robust ResNet50 model, and activations are extracted from the conv1, layer3, and avgpool layers. For each layer, filter-by-filter Gram matrices are computed and vectorized. Using these vectorized features from the ImageNet dataset, ICA is applied, and the top three components with the highest explained variance are selected as exploration axes for each layer.
@@
A Gram matrix represents the similarity of activations across multiple filters, summed over spatial positions. Given the activation of the $l$-th layer of a CNN for an input image as $F_{jk}$, where $j$ denotes the filter index and $k$ denotes the spatial position, the Gram matrix $G_{ij}$ is defined as:
$G_{ij}=\sum_{k}F_{ik}F_{jk}.$ (1)
@@
In this study, Gram matrices were extracted from multiple layers of the model using 1,000 images randomly selected from the ImageNet validation dataset (Deng et al.,, 2009).
After extracting the Gram matrix features, ICA was performed on them, and components with high explanatory power were selected, thereby considering images as points in the ICA coordinate space. The exploration direction was defined along the axes of ICA from the reference stimulus.
Although several dimensionality reduction methods could be considered, we used ICA for the following reasons. First, principal component analysis (PCA), which preserves orthogonality in the input space, is not suitable for disentangling high-level representations in DNNs (as described in Fig. S2). In addition, ICA standardizes components to have unit variance, making it easier to compare representations across different layers.
@@
In the following, we describe in detail how dimensionality reduction of the metameric space was performed using ICA. ICA assumes non-Gaussianity of signals and computes a linear transformation $\mathbf{W}$ to separate observed signals $\mathbf{X}$ into independent components $\mathbf{S}$ as $\mathbf{S}=\mathbf{W}\mathbf{X}$. Each independent component (hereafter referred to as an ICA component) is standardized to have zero mean and a covariance matrix equal to the identity matrix. The Gram matrices of the target layers were vectorized to form feature vectors $\mathbf{x}$ for each image. Using the feature vector $\mathbf{x}$ of 1,000 images, an input data matrix $\mathbf{X}$ was constructed, and ICA was applied to obtain the independent component matrix $\mathbf{S}$:
$\mathbf{S}=\mathbf{W}\mathbf{X},\quad\mathbf{s}=\mathbf{W}\mathbf{x},\quad\mathbf{X}=\mathbf{S}\mathbf{A}^{\mathsf{T}}$ (2)
@@
where $A$ represents the mixing matrix. The number of independent components was pre-set to 100, and the top 3 components with the highest contribution were selected. Practically, we used FastICA of the Python scikit-learn library to compute the components. The contribution of each component was quantified using the explained variance (EV) defined as:
$\mathrm{EV}_{i}=1-\frac{\|\mathbf{X}-\hat{\mathbf{X}}_{i}\|_{\text{F}}^{2}}{\|\mathbf{X}\|_{\text{F}}^{2}},$ (3)
@@
$\hat{\mathbf{X}}_{i}=\mathbf{S_{i}}\mathbf{A}^{\mathsf{T}},$ (4)
with $\mathbf{S_{i}}$ being the matrix retaining only the $i$-th independent component. Unlike PCA, the total explained variance for multiple ICA components does not necessarily match the sum of the individual component contributions. The contribution of each component is shown in Fig. S4.
Finally, we computed the metameric space based on features extracted from the conv1, layer3, and avgpool layers of the robust ResNet50. These layers were selected to examine early, intermediate, and higher-level stages of visual processing, while avoiding layers affected by skip connections in the ResNet architecture. Both positive and negative directions of each ICA component were used in the exploration.
2.3 Generation of Target Images
This section describes the method for generating images at designated positions using gradient descent. The image modification process was performed by specifying the target layer (conv1, layer3, and avgpool), the ICA component of interest (components 0, 1, 2), the direction of change (positive or negative), the magnitude of change (a positive scalar corresponding to the target value), and the reference image (i.e., the initial image of the modification).
The image generation procedure is as follows:
  1. 1.
    Compute the ICA components of the reference image for the target layer as $\mathbf{y}_{\text{o}}=\mathbf{W}\mathbf{x}_{\text{o}}$, where $\mathbf{x}_{\text{o}}$ is the vectorized Gram matrix of the target layer.
  2. 2.
    Define the target component as $\mathrm{p}$ and the target value as $t$. The modified ICA component $\mathbf{y}_{\text{t}}$ is given by:
    $\mathbf{y}_{\text{t}}=\mathbf{y}_{\text{o}}+t\mathbf{e}_{\text{p}},$ (5)
    where $\mathbf{e}_{i}$ is a unit vector with 1 at the $i$-th component and 0 elsewhere. This operation modifies only the target component while keeping all others unchanged. If the change direction is negative, the unit vector is multiplied by 1-1.
  3. 3.
    In each iteration, compute the ICA components of the current image as $\mathbf{y}_{\text{c}}=\mathbf{W}\mathbf{x}_{\text{c}}$.
  4. 4.
    Compute the mean squared error (MSE) between $\mathbf{y}_{\text{t}}$ and $\mathbf{y}_{\text{c}}$ as the loss function.
  5. 5.
Examples of image generation for the three layers and three components used in the experiment are shown in Fig.4 and Fig. S3.
Figure 4: An example of the images used in the experiment, generated with three layers and three components each. The numbers (in red and blue) below the layer names represent the target ICA values used during image generation.The magnitude of modulation in this visualization is scaled according to the experimental results.
The selection of reference images for initialization was based on two criteria. First, we aimed to use images located near the zero point of the ICA components. Second, to prevent participants in the behavioral experiment from memorizing individual images, we needed to sample many images. For this purpose, from 50,000 images in the ImageNet validation dataset, we sampled images for three target layers whose ICA component distances from the origin were within the lowest 20%. Among these, 614 images that met the selection criteria across all three layers were used as experimental stimuli (Table 1).